OpenStudy (anonymous):

Someone please help me on this, I asked this over 30 minutes ago. Give the slope of each line then determine whether the two lines are parallel, perpendicular, or neither parallel or perpendicular. 5x+4y=0 -8x+10y=0

OpenStudy (kropot72):

First put each equation in the form y = mx + b. The first equation then becomes y = -(5/4)x + 0 Can you carry out this for the second given equation -8x + 10y = 0 ?

OpenStudy (anonymous):

y=(8/10)x+0?

OpenStudy (kropot72):

Rearranging the second given equation -8x + 10y = 0 into the form y = mx + b is done as follows: Adding 8x to both sides we get 10y = 8x. Dividing both sides by 10 gives y = (8/10)x + 0 Do you know what m is in the equation y = mx + b ?

OpenStudy (anonymous):

The slope right?

OpenStudy (kropot72):

Yes, m stands for the slope. So from the previous work can you find the slope of each line?

OpenStudy (anonymous):

The slope would be the 8/10 and the 5/4

OpenStudy (kropot72):

Nearly right!. The slopes are 8/10 and -(5/4). Note that 8/10 = 4/5. If the lines were parallel the slopes would be the same. The lines are not parallel because the slopes are different. However the slopes are the negative inverse of each other. So can we say the lines are perpendicular?

OpenStudy (anonymous):

Yes.

OpenStudy (kropot72):

That is correct. The lines are perpendicular.